OpenStudy (anonymous):

Evaluate the following expression for x = –2 and y = –4. 4(x-y)^2

OpenStudy (anonymous):

4(-2 - -4)^2

jigglypuff314 (jigglypuff314):

yes so what does -2 - (-4) = ?

OpenStudy (anonymous):

-6

OpenStudy (shamil98):

4(-2-(-4)^2 4(2)^2 4(4) =your answer..

jigglypuff314 (jigglypuff314):

@shamil98 could you help explain how -2 - (-4) = 2 and not -6

OpenStudy (shamil98):

- - = + + - = - ++ = + - + = -

OpenStudy (shamil98):

-2-(-4) -2+ 4 =2

OpenStudy (shamil98):

(x-y) y = -4 x-(-4) ......